# Distribution Prompting: Understanding the Expressivity of Language Models Through the Next-Token Distributions They Can Produce

### v1

@@ Abstract
Haojin Wang1 Zining Zhu 2 Freda Shi 1, 3 1 University of Waterloo 2 Stevens Institute of Technology 3 Vector Institute, Canada CIFAR AI Chair {s2323wang,fhs}@uwaterloo.cazzhu41@stevens.edu
@@ 1 Introduction
To answer this question, we conduct a systematic analysis of the expressiveness of LMs, in terms of approximating different target probability distributions. Without changing the parameters of LMs, we aim to find prompts that lead to the output of a probability distribution as close as possible to a target distribution. Apparently, due to inefficiency for enumerating the infinitely many prompts as triggers for approximating the target distribution, we use modern optimization techniques. Through either soft (Li and Liang, 2021) or hard (Wallace et al., 2019) [^1] The key difference between soft and hard prompts is that soft prompts can be continuous embeddings that do not align to any token in the vocabulary, while hard prompts must correspond to vocabulary entries. prompt tuning across different initializations, we measure the difficulty of an approximation using the minimum Kullback–Leibler (KL) divergence between the target and the approximated distributions from different prompt initializations.
We conduct experiments on multiple pretrained LMs, including GPT-2 (Radford et al., 2019), Llama3.2-1B (Dubey et al., 2024), Falcon-7B (Almazrouei et al., 2023), and Llama3-8B (Dubey et al., 2024). We identify that these models share general properties for the next-token probability distributions: First, as illustrated in Fig. 1, the difficulty of approximating a random target distribution is approximately a unimodal function of the target entropy, where the approximation loss increases with the entropy of the target distribution up to a point, then decreases across all models. Second, for the distributions with moderate entropy, the existence of outlier tokens (i.e., those receiving much higher probability values than the other tokens) makes the approximation easier. Finally and unsurprisingly, LM-generated distributions are easier to approximate regardless of the entropy values, and can be effectively transferred across LMs. Our results introduce novel evidence toward a better understanding of the expressiveness of LMs, highlighting the opportunities and challenges in using LMs as probability distribution proposers.
@@ 2 Related Work
Figure 2: Illustration of the prompt tuning frameworks adapted in this work. For hard (Fig. 2(a)) and hybrid (Fig. 2(c)) embeddings–based frameworks, we optimize the prefix embeddings from left to right; forthe soft embeddings-based framework (Fig. 2(b)), we optimize the embeddings directly through gradient descent. Best viewed in color: the dark colors represent the initial embeddings, and the light colors represent the optimized ones.
@@ 3 Problem Formulation
Consider a Transformer-based (Vaswani et al., 2017) autoregressive language model $\mathcal{M}_{\boldsymbol{\Phi}}$ (e.g., Llama3; Dubey et al., 2024) parametrized by ${\boldsymbol{\Phi}}$. Let $\mathbf{x}=[x_{1},x_{2},...,x_{n}]\in\mathbb{Z}_{+}^{n}$ be the input text with $n$ tokens. The language model $\mathcal{M}_{\boldsymbol{\Phi}}$ first converts $\mathbf{x}$ into non-contextualized token embeddings $\mathbf{W}_{\boldsymbol{\Phi}}(\mathbf{x})=[\mathbf{w}_{\boldsymbol{\Phi}}(x_{1}),\ldots,\mathbf{w}_{\boldsymbol{\Phi}}(x_{n})]\in\mathbb{R}^{n\times d}$ through a look-up table, where $d$ is the embedding size. The embeddings are then fed into the Transformer to produce the logits over the vocabulary $\mathbf{z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{n}]\in\mathbb{R}^{|\mathcal{V}|}$, where $|\mathcal{V}|$ is the size of the vocabulary $\mathcal{V}$. The logits are later on normalized with a softmax operator to produce the probability distribution $\mathbf{q}$ over the vocabulary, where for $v\in\mathcal{V}$, the probability of the next token being $v$ given the prefix $\mathbf{x}$ is given by: [^2] With a slight abuse of notations, we use $v$ to denote both the token and the token index in the vocabulary.
@@
We aim to find an input sequence $\mathbf{x}$, where $\mathbf{q}=\mathcal{M}_{\boldsymbol{\Phi}}(\mathbf{x})$ is close enough to an arbitrary target distribution $\mathbf{p}\in\mathbb{R}^{|\mathcal{V}|}$. Here, since we adopt optimization techniques (Li and Liang, 2021; Wallace et al., 2019) to find $\mathbf{x}$, $\mathbf{x}$ can be viewed as parameters $\boldsymbol{\Theta}$ ($\boldsymbol{\Theta}=\mathbf{x}$) optimized by the algorithm (Fig. 2). The objective of the optimization problem is to minimize the KL divergence between the target distribution $\mathbf{p}$ and the output distribution $\mathbf{q}$:
@@
Lower KL-divergence indicates better approximation of the target distributions. Varying across random seeds, the minimum KL-divergence loss $\mathcal{L}_{\text{KL}}(\mathbf{p},\mathbf{q};{\boldsymbol{\Theta}})$ gives an approximation on how difficult it is to elicit the target distribution $\mathbf{p}$ from the LM $\mathcal{M}_{\boldsymbol{\Phi}}$—lower loss indicates easier elicitation.
@@ 4 Methods
In this section, we introduce our method to synthesize target probability distributions (§ 4.1) and the prompt tuning frameworks (§ 4.2).
@@ 4.1 Target Distribution Construction
$\mathcal{L}(\mathbf{z};h)$ $=||H(p(\cdot;\mathbf{z}))-h||^{2}$ (1)
$=\left|\left|-\left(\sum_{v\in\mathcal{V}}p(v;\mathbf{z})\log p(v;\mathbf{z})\right)-h\right|\right|^{2},$
@@
However, entropy is a lossy descriptor of a probability distribution. For example, in Fig. 3, the two distributions (one output by Llama3.2-1B (Dubey et al., 2024) and the other by our search algorithm based on Eq. 1) have the same entropy yet are distinct from each other sharply. Considering the fact the LMs usually generate skewed distribution with a few tokens receiving much higher probability mass than the rest (Holtzman et al., 2020), we define distributions with outliers, of which a large amount of probability mass concentrates on a small set of tokens. Formally, a distribution with $k$ outliers satisfies:
@@
where $\mathcal{V}_{o}$ denotes the $k(k<<|\mathcal{V}|)$ pre-selected outlier tokens that receive distinctively larger probability mass than the rest; $m$ is the theoretical upper bound of the concentrated probability mass to $k$ tokens to ensure the feasibility of the entire distribution reaching the target entropy $h$—details of calculating $m$ from $k$, $h$, and $|\mathcal{V}|$ can be found in App. A; $\delta$ is a small margin to ensure the feasibility of the optimization process, which we set to $0.01$ in all experiments.
We add a regularization term to the loss function in Eq. 1 to encourage increased probability mass on the outlier tokens. To construct an outlier distribution, we optimize the following loss function with gradient descent:
@@
where $\alpha$ and $\beta$ are hyperparameters that balance the two terms, and stop once Eq. 2 is satisfied. We will compare the difficulty of approximating the distributions generated by Eq. 1 and Eq. 3 in our experiments.
Figure 3: Two distributions, both with entropy 7.68 but have distinctively different shapes. Upper: the output probability distribution from prompting Llama3.2-1B with ’I love’. Lower: The distribution from our search algorithm following the objective of Eq. 1.
@@ 4.2 Prompt Tuning Frameworks
In practice, we use AdamW (Loshchilov and Hutter, 2019) to optimize this objective.
Hybrid prompt tuning. To better understand the effect of different tuning methods, we introduce a hybrid tuning method that combines hard and soft prompts, where $\boldsymbol{\Theta}$ can be partitioned into two parts: $\boldsymbol{\Theta}_{h}$ for hard prompts and $\boldsymbol{\Theta}_{s}$ for soft prompts, with some $\boldsymbol{\theta}$ updated by hard prompts tuning and others by soft prompts tuning. We also optimize the hybrid embeddings in a left-to-right manner, where the hard embeddings are updated using Eq. 4, with soft embeddings updated using Eq. 5.
We compare the above three strategies in approximating target distributions. The general framework of our tuning methods can be found in Fig. 2.
@@ 5 Experiments
To study the general property of the expressiveness of an LM, we start by studying different probability distributions of different entropy (§ 5.1). We then analyze the LM approximations performance on outlier distributions (§ 5.2). Finally, we investigate target probability distributions as the output distribution of LM itself (§ 5.3), and analyze the results for approximating the variations of such distributions (§ 5.4). We set up our experiments across transformer-based LMs of different architectures and sizes, including GPT-2 (Radford et al., 2019), Llama3.2-1B (Dubey et al., 2024), Falcon-7B (Almazrouei et al., 2023), and Llama3-8B (Dubey et al., 2024). To mitigate the effect of randomness brought by target distribution construction algorithms in § 4.1, we obtain different distributions regarding the same entropy value. For each distribution, we use a different initialization for the prefix parameters. Detailed experiment setups can be found in App. B for reference.
@@ 5.1 Approximation of Vanilla Distributions
We perform prompt tuning to approximate target distributions constructed with Eq. 1, which we will refer to as vanilla distributions. We investigate the target entropy ranging from 0 to $\log|\mathcal{V}|$ with a step size 0.05, where $|\mathcal{V}|$ is the model’s vocabulary size, and $\log|\mathcal{V}|$ is the theoretical upper bound for the vocabulary distribution. [^3] The proof is detailed in App. A.
Fig. 1 presents the approximation loss of soft prompt tuning, plotted against the target entropy for different models. The approximation loss curve exhibits a consistent pattern across all models: increasing with target entropy up to a point, then decreasing across all models, which roughly forms a unimodal curve.
@@
We then compare the soft and hard prompt tuning frameworks (§ 4.2) on the vanilla target distributions (Fig. 4). As desired, since soft prompts are not constrained to represent actual vocabulary tokens, they generally achieve lower approximation losses than hard prompts. To that end, we incorporate hybrid prompt tuning that executes updates based on the embedding source, which helps understand the gap between hard and soft prompt tuning. As we increase the number of soft tokens in the prefix, approximation losses improved, demonstrating the advantage of soft tuning. In addition, the approximation loss of different random initializations of the prefix embeddings may lead to significantly different results. We find this phenomenon is particularly pronounced under hybrid prompt tuning (Fig. 4(b)) and soft prompt tuning (Fig. 4(c))—loosely connected with Zhao et al. (2025), our results also support that randomness plays a systematic role in minimizing the loss.
@@
Figure 6: Approximation losses exhibit a consistent pattern as shown in Fig. 1 for models randomly initialized who excluded pretraining history. The orange-yellow line is the randomly initiated models, while the blue dash line corresponds to the pretrained models. The x-axis in the subplot stands for target entropy.
However, even when using pure soft prompts for these objectives, we still struggled to elicit certain distributions. We present the result of a soft prompt tuning study on Llama3.2-1B in Fig. 5. The result shows systematic trends of target distributions regarding different ranges of entropy values. At low entropy, we have near-zero losses on soft prompt approximation, and the prompts tuned by our framework output distribution with aligned entropy to the target. However, when the target distribution has a larger entropy, the losses of our training objective increase due to the failure of the model to approximate a lower-entropy distribution from the initializations—the entropy of the approximated distribution grows to be larger than the target entropy. Such phenomenon in approximation loss is also observed for transformer-based models from different series and sizes (Fig. 1). When the parameters are randomly initialized, the approximation loss remains a unimodal curve with respect to the target entropy (Fig. 6), indicating that pretraining is not the cause of the unimodal difficulty phenomenon.
@@
Figure 8: The comparison of approximation performance across distribution pairs in § 5.3 on Llama3.2-1B.
@@ 5.2 Approximation of Outlier Distributions
In this section, we perform prompt tuning to approximate target distributions characterized by outlier tokens, where these distributions are constructed according to the objective defined in Eq. 3.
@@
Given that distributions with moderate entropy exhibits higher approximation losses in the vanilla setting (Fig. 5), we focus our investigation on this range of entropy values, where the approximation difficulty is most pronounced. Our results (Figs. 7(a) and 7(b)) demonstrate that target distributions containing outlier tokens achieve lower approximation losses compared to vanilla distributions with comparable entropy levels. This improvement is particularly pronounced when a single outlier token receives a substantially higher probability mass relative to the remaining tokens. In these cases, the tuned prompts also yield output distributions with entropy values more closely aligned to those of the target distributions. These findings suggest that the presence of outliers in the target distribution may facilitate more effective prompt tuning by the LM.

### v2

@@ Abstract
Haojin Wang1thanks: Now at the University of Illinois Urbana-Champaign. Work done during visiting term at the University of Waterloo.   Zining Zhu 2  Freda Shi 1, 3 1 University of Waterloo 2 Stevens Institute of Technology 3 Vector Institute, Canada CIFAR AI Chair {s2323wang,fhs}@uwaterloo.cazzhu41@stevens.edu
@@ 1 Introduction
To answer this question, we conduct a systematic analysis of the expressiveness of LMs, in terms of approximating different target probability distributions. Without changing the parameters of LMs, we aim to find prompts that lead to the output of a probability distribution as close as possible to a target distribution. This design isolates the intrinsic expressivity of pretrained LMs, since we probe their behavior by varying only the prompts while keeping model parameters fixed. Apparently, due to inefficiency for enumerating the infinitely many prompts as triggers for approximating the target distribution, we use modern optimization techniques. Through either soft (Li and Liang, 2021) or hard (Wallace et al., 2019) [^1] The key difference between soft and hard prompts is that soft prompts can be continuous embeddings that do not align to any token in the vocabulary, while hard prompts must correspond to vocabulary entries. prompt tuning across different initializations, we measure the difficulty of an approximation using the minimum Kullback–Leibler (KL) divergence between the target and the approximated distributions from different prompt initializations.
We conduct experiments on multiple pretrained LMs, including GPT-2 (Radford et al., 2019), Llama3.2-1B (Dubey et al., 2024), Falcon-7B (Almazrouei et al., 2023), and Llama3-8B (Dubey et al., 2024). We identify that these models share general properties for the next-token probability distributions: First, as illustrated in Fig.˜1, the difficulty of approximating a random target distribution is approximately a unimodal function of the target entropy, where the approximation loss increases with the entropy of the target distribution up to a point, then decreases across all models. Second, for the distributions with moderate entropy, the existence of outlier tokens (i.e., those receiving much higher probability values than the other tokens) makes the approximation easier. Finally and unsurprisingly, LM-generated distributions are easier to approximate regardless of the entropy values, and can be effectively transferred across LMs. Our results introduce novel evidence toward a better understanding of the expressiveness of LMs, highlighting the opportunities and challenges in using LMs as probability distribution proposers.
@@ 2 Related Work
Figure 2: Illustration of the prompt tuning frameworks adapted in this work. For hard (Fig.˜2(a)) and hybrid (Fig.˜2(c)) embeddings–based frameworks, we optimize the prefix embeddings from left to right; forthe soft embeddings-based framework (Fig.˜2(b)), we optimize the embeddings directly through gradient descent. Best viewed in color: the dark colors represent the initial embeddings, and the light colors represent the optimized ones.
@@ 3 Problem Formulation
Consider a Transformer-based (Vaswani et al., 2017) autoregressive language model $\mathcal{M}_{\boldsymbol{\Phi}}$ (e.g., Llama3; Dubey et al., 2024) parametrized by $\boldsymbol{\Phi}$}. Let $\mathbf{x}=[x_{1},x_{2},...,x_{n}]\in\mathbb{Z}_{\geq 0}^{n}$ be the input text with $n$ tokens. The language model $\mathcal{M}_{\boldsymbol{\Phi}}$ first converts $\mathbf{x}$ into non-contextualized token embeddings $\mathbf{W}_{\boldsymbol{\Phi}}(\mathbf{x})=[\mathbf{w}_{\boldsymbol{\Phi}}(x_{1}),\ldots,\mathbf{w}_{\boldsymbol{\Phi}}(x_{n})]\in\mathbb{R}^{n\times d}$ through a look-up table, where $d$ is the embedding size. The embeddings are then fed into the Transformer to produce the logits over the vocabulary $\mathbf{z}=[z_{1},\ldots,z_{n}]\in\mathbb{R}^{|\mathcal{V}|}$, where $|\mathcal{V}|$ is the size of the vocabulary $\mathcal{V}$. The logits are later on normalized with a softmax operator to produce the probability distribution $\mathbf{q}$ over the vocabulary, where for $v\in\mathcal{V}$, the probability of the next token being $v$ given the prefix $\mathbf{x}$ is given by: [^2] With a slight abuse of notations, we use $v$ to denote both the token and the token index in the vocabulary.
@@
We aim to find an input sequence $\mathbf{x}$, where $\mathbf{q}=\mathcal{M}_{\boldsymbol{\Phi}}(\mathbf{x})$ is close enough to an arbitrary target distribution $\mathbf{p}\in\mathbb{R}^{|\mathcal{V}|}$. Here, since we adopt optimization techniques (Li and Liang, 2021; Wallace et al., 2019) to find $\mathbf{x}$, $\mathbf{x}$ can be viewed as parameters $\boldsymbol{\Theta}$ ($\boldsymbol{\Theta}=\mathbf{x}$) optimized by the algorithm (Fig.˜2). The objective of the optimization problem is to minimize the KL divergence between the target distribution $\mathbf{p}$ and the output distribution $\mathbf{q}$:
@@
Lower KL divergence indicates better approximation of the target distributions. Varying across random seeds, the minimum KL-divergence loss $\mathcal{L}_{\text{KL}}(\mathbf{p},\mathbf{q};{\boldsymbol{\Theta}})$ gives an approximation on how difficult it is to elicit the target distribution $\mathbf{p}$ from the LM $\mathcal{M}_{\boldsymbol{\Phi}}$—lower loss indicates easier elicitation.
@@ 4 Methods
In this section, we introduce our method to synthesize target probability distributions (§˜4.1) and the prompt tuning frameworks (§˜4.2).
@@ 4.1 Target Distribution Construction
[…]
@@
However, entropy is a lossy descriptor of a probability distribution. For example, in Fig.˜3, the two distributions (one output by Llama3.2-1B (Dubey et al., 2024) and the other by our search algorithm based on Eq.˜1) have the same entropy yet are distinct from each other sharply. Considering the fact that LMs usually generate skewed distributions with a few tokens receiving much higher probability mass than the rest (Holtzman et al., 2020), we define distributions with outliers, of which a large amount of probability mass concentrates on a small set of tokens. Formally, a distribution with $k$ outliers satisfies:
@@
where $\mathcal{V}_{o}$ denotes the $k(k\ll|\mathcal{V}|)$ pre-selected outlier tokens that receive distinctively larger probability mass than the rest; $m$ is the theoretical upper bound of the concentrated probability mass to $k$ tokens to ensure the feasibility of the entire distribution reaching the target entropy hh—details of calculating $m$ from $k$, $h$, and $|\mathcal{V}|$ can be found in App.˜A; $\delta$ is a small margin to ensure the feasibility of the optimization process, which we set to $0.01$ in all experiments. Intuitively, Eq.˜2 requires that a small set of $k$ tokens account for most of the probability mass, thereby capturing our notion of outlier distributions. Here, $m$ specifies the maximum feasible cumulative mass these tokens can hold under the entropy constraint, ensuring the construction remains mathematically valid.
We add a regularization term to the loss function in Eq.˜1 to encourage increased probability mass on the outlier tokens. To construct an outlier distribution, we optimize the following loss function with gradient descent:
@@
where $\alpha$ and $\beta$ are hyperparameters that balance the two terms, and stop once Eq.˜2 is satisfied. We will compare the difficulty of approximating the distributions generated by Eq.˜1 and Eq.˜3 in our experiments.
Figure 3: Two distributions, both with entropy 7.68 but have distinctively different shapes. Upper: the output probability distribution from prompting Llama3.2-1B with ’I love’. Lower: The distribution from our search algorithm following the objective of Eq.˜1.
@@ 4.2 Prompt Tuning Frameworks
We use AdamW (Loshchilov and Hutter, 2019) to optimize this objective. We include soft prompts to probe the upper bound of LM expressivity, since they allow unconstrained optimization in the embedding space; although they may not lie on the natural input manifold, we aim to study theoretical reachability rather than mimic typical usage.
Hybrid prompt tuning. To better understand the effect of different tuning methods, we introduce a hybrid tuning method that combines hard and soft prompts, where $\boldsymbol{\Theta}$ can be partitioned into two parts: $\boldsymbol{\Theta}_{h}$ for hard prompts and $\boldsymbol{\Theta}_{s}$ for soft prompts, with some $\boldsymbol{\theta}$ updated by hard prompts tuning and others by soft prompts tuning. We also optimize the hybrid embeddings in a left-to-right manner, where the hard embeddings are updated using Eq.˜4, with soft embeddings updated using Eq.˜5.
We compare the above three strategies in approximating target distributions. The general framework of our tuning methods can be found in Fig.˜2.
@@ 5 Experiments
To study the general property of the expressiveness of an LM, we start by studying different probability distributions of different entropy (§˜5.1). We then analyze the LM approximation performance on outlier distributions (§˜5.2). Finally, we investigate target probability distributions as the output distribution of LM itself (§˜5.3), and analyze the results for approximating the variations of such distributions (§˜5.4). We set up our experiments across transformer-based LMs of different architectures and sizes, including GPT-2 (Radford et al., 2019), Llama3.2-1B (Dubey et al., 2024), Falcon-7B (Almazrouei et al., 2023), and Llama3-8B (Dubey et al., 2024). To mitigate the effect of randomness brought by target distribution construction algorithms in §˜4.1, we obtain different distributions regarding the same entropy value. For each distribution, we use a different initialization for the prefix parameters. Detailed experiment setups can be found in App.˜B for reference.
@@ 5.1 Approximation of Vanilla Distributions
We perform prompt tuning to approximate target distributions constructed with Eq.˜1, which we will refer to as vanilla distributions. We investigate the target entropy ranging from 0 to $\log|\mathcal{V}|$ with a step size 0.05, where $|\mathcal{V}|$ is the model’s vocabulary size, and $\log|\mathcal{V}|$ is the theoretical upper bound for the vocabulary distribution. [^3] The proof is detailed in App. A.
Fig.˜1 presents the approximation loss of soft prompt tuning, plotted against the target entropy for different models. The approximation loss curve exhibits a consistent pattern across all models: increasing with target entropy up to a point, then decreasing across all models, which roughly forms a unimodal curve.
@@
We then compare soft and hard prompt tuning frameworks (§˜4.2) on vanilla target distributions (Fig.˜4). As expected, soft prompts—being unconstrained by the vocabulary—consistently yield lower approximation losses than hard prompts. To further explore this gap, we adopt hybrid prompt tuning, where updates depend on the embedding source. Increasing the number of soft tokens in the prefix improves approximation, highlighting the benefit of soft tuning. We also find that different random initializations of prefix embeddings can lead to large variance in performance. This effect is especially pronounced in hybrid and soft prompt tuning (Figs.˜4(b) and 4(c)). Echoing findings in Zhao et al. (2025), our results suggest that randomness plays a systematic role in loss minimization.
@@
Figure 6: Approximation losses exhibit a consistent pattern as shown in Fig.˜1 for models randomly initialized that excluded pretraining history. The orange-yellow line is the randomly initiated models, while the blue dash line corresponds to the pretrained models. The x-axis in the subplot stands for target entropy.
However, even when using pure soft prompts for these objectives, we still struggled to elicit certain distributions. We present the result of a soft prompt tuning study on Llama3.2-1B in Fig.˜5. The result shows systematic trends of target distributions regarding different ranges of entropy values. At low entropy, we have near-zero losses on soft prompt approximation, and the prompts tuned by our framework output a distribution with aligned entropy to the target. However, when the target distribution has a larger entropy, the losses of our training objective increase due to the failure of the model to approximate a lower-entropy distribution from the initializations—the entropy of the approximated distribution grows to be larger than the target entropy. Such a phenomenon in approximation loss is also observed for transformer-based models from different series and sizes (Fig.˜1). When the parameters are randomly initialized, the approximation loss remains a unimodal curve with respect to the target entropy (Fig.˜6), indicating that pretraining is not the cause of the unimodal difficulty phenomenon.
@@
Figure 8: The comparison of approximation performance across distribution pairs in §˜5.3 on Llama3.2-1B.
@@ 5.2 Approximation of Outlier Distributions
In this section, we perform prompt tuning to approximate target distributions characterized by outlier tokens, where these distributions are constructed according to the objective defined in Eq.˜3.
@@
Since moderate-entropy distributions yield higher approximation losses in the vanilla setting (Fig.˜5), we focus our analysis on this regime, where approximation is most challenging. Our results (Figs.˜7(a) and 7(b)) show that distributions with outlier tokens incur lower losses than vanilla counterparts with similar entropy. This effect is strongest when a single token dominates the probability mass. In such cases, the tuned prompts also produce output distributions whose entropy better matches the target. These findings suggest that the presence of outliers in the target distribution may facilitate more effective prompt tuning by the LM.